OpenStudy (anonymous):

Find the derivative of f(x) = 3x + 8 at x = 4. Help please :(

OpenStudy (abb0t):

derivative is 3

OpenStudy (anonymous):

thanks but how did you get it?

OpenStudy (anonymous):

first differentiate the equation this would be f'(x)=3+0 now put value of x and get the result which is 3

OpenStudy (anonymous):

im confused, i dont know how you got that :(

OpenStudy (anonymous):

basic rule of differentiation! if u differentiate anything of the form nx where n is a constant you get the constant after differentiating. and if u differentiate a constant like 8 you get 0

OpenStudy (anonymous):

use the power rule which is taking the power moving it down and subtracting it by 1 so \[\frac{d}{dx}x^{n}=nx^{n-1}\]

OpenStudy (abb0t):

\(\frac{d}{dx}a^n = na^{n-1}\) in this case n = 1 and 1-1 = 0 and \(x^0=1\)

OpenStudy (anonymous):

im still confused guys, im sorry :/

OpenStudy (anonymous):

@britanny_ and how is that possible??

OpenStudy (abb0t):

Britanny, what are you doing? Go home, you're drunk.

OpenStudy (anonymous):

@britanny_ u need the derivative first i think! if you put the value then then you wont get the answer!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok guys o-o be nice

OpenStudy (anonymous):

that kinda just made my day (sadly )

OpenStudy (anonymous):

fine whatever you answer it in a way she can understand.

OpenStudy (anonymous):

:/ could you help me @britanny_ i have like 4 more of these to do and im confused :(

OpenStudy (agent0smith):

@britanny_ i think you missed that this is a calculus question - your method is not correct here. @iluvvyyhu do you know how to find the derivative of y = 3x +8?

OpenStudy (anonymous):

no i dont :/ and the class im taking is precalc but its online and they didnt have a good explanation

OpenStudy (anonymous):